OpenStudy (anonymous):

volume of a cone solve for h V=(1/3)pie r^2h

OpenStudy (anonymous):

this is just like the other ones.. just divide both sides by the 'non h' factors

OpenStudy (anonymous):

would it be like this?: v=91/3)piR^2h 3v=piR^2h 3v/pi-R^2h (3v)/pi(R^2)=h h=(3v)/(pi)R^2